OpenStudy (anonymous):

if 6y+3x+8xy-94=0 and x+5=13 then y+4=? plz show work :D

jimthompson5910 (jim_thompson5910):

x+5=13 x = ??

OpenStudy (anonymous):

x=8

jimthompson5910 (jim_thompson5910):

6y+3x+8xy-94=0 6y+3*8+8*8y-94=0 ... plug in x = 8 6y+24+64y-94=0 70y-70 = 0 y = ??

OpenStudy (anonymous):

oh 1....

OpenStudy (anonymous):

thankk you!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

so if y = 1, then y+4 = 1+4 = 5

jimthompson5910 (jim_thompson5910):

np